OpenStudy (anonymous):

PLEASE HELP!!! MEDAL WILL BE REWARDED! WHAT RULE DO I USE? FIND Y': y=3x^2+^3(sqrt x^4)

jimthompson5910 (jim_thompson5910):

You use the rule \[\large \frac{d}{dx}(x^n) = nx^{n-1}\] and the idea that \[\large \sqrt[n]{x^m} = x^{m/n}\]

OpenStudy (anonymous):

So is that mean I would use product rule?

jimthompson5910 (jim_thompson5910):

no because \[\Large y = 3x^2 + \sqrt[3]{x^4}\] is a sum (not a product)

OpenStudy (anonymous):

oh so quotient?

jimthompson5910 (jim_thompson5910):

you use the sum rule \[\large \frac{d}{dx}[f(x) + g(x)] = \frac{d}{dx}[f(x)] + \frac{d}{dx}[g(x)]\]

jimthompson5910 (jim_thompson5910):

basically: the derivative of the sum of two expressions/functions is the same as deriving the two functions individually first, then adding

OpenStudy (anonymous):

oh okay, so I wold get dy dx =9x^2 x √ 3 +x 7/3

jimthompson5910 (jim_thompson5910):

hold on, is the equation \[\Large y = 3x^2 + \sqrt[3]{x^4}\] OR is it \[\Large y = 3x^2 * \sqrt[3]{x^4}\]

OpenStudy (anonymous):

its +

jimthompson5910 (jim_thompson5910):

ok, so you would do it like this \[\Large y = 3x^2 + \sqrt[3]{x^4}\] \[\Large y = 3x^2 + x^{4/3}\] \[\Large \frac{d}{dx}[y] = \frac{d}{dx}[3x^2 + x^{4/3}]\] \[\Large \frac{d}{dx}[y] = \frac{d}{dx}[3x^2] + \frac{d}{dx}[x^{4/3}]\] \[\Large \frac{d}{dx}[y] = 3*2x^{2-1} + \frac{4}{3}x^{4/3-1}\] \[\Large \frac{d}{dx}[y] = 3*2x^{1} + \frac{4}{3}x^{1/3}\] \[\Large \frac{d}{dx}[y] = 6x + \frac{4}{3}x^{1/3}\] \[\Large y^{\prime} = 6x + \frac{4}{3}x^{1/3}\]

OpenStudy (anonymous):

ohhhh okay, im sorry I was using dx/dy instead of d/dx

jimthompson5910 (jim_thompson5910):

you would use either dy/dx, f'(x), y' or d/dx[ ... ] (where you replace the dot dot dot with the expression) dx/dy is going the wrong way

OpenStudy (anonymous):

Oh okay, omg thank you very much!!!

jimthompson5910 (jim_thompson5910):

you're welcome